# Assessing Small Language Models for Code Generation: An Empirical Study with Benchmarks

### v3

@@ 4.2 Efficiency Trade-offs in SLM Deployment (RQ2)
Resource Consumption and Inference Speed vs. Performance of SLMs: Our analysis reveals that GPU memory usage (VRAM) increases significantly with model size. At the same time, inference speed does not show a corresponding statistical variation. This indicates that, rather than inference speed, VRAM is the constraint when deploying SLMs in practice. As models scale up, their memory demands grow higher, creating significant challenges for deployment in resource-constrained environments. However, this does not imply that larger models are always necessary to achieve high performance. While high-performing SLMs often belong to the larger parameter groups, several smaller and mid-sized models demonstrate a favourable balance between correctness and resource efficiency. These models offer compelling alternatives for use cases where hardware limitations prohibit the use of resource-intensive models.
Prior studies (e.g., [8], [5]) have similarly emphasized this dynamic. While inference speed can often be optimised through software and hardware tuning, memory capacity, especially VRAM, remains a primary constraint in deploying small language models. Recent findings by Ashkboos et al. [21] indicate that even SLMs with modest parameter counts face significant memory bottlenecks during both training and inference, limiting their usability in resource-constrained settings. Lu et al. [8] also observed that many performance evaluations overlook memory bottlenecks, leading to unrealistic assessments of deployment feasibility. Moreover, Subramanian et al. [3] emphasised treating efficiency not merely as another dimension of scale, but as an important metric in empirical evaluations of SLMs, especially in applied software engineering contexts. Furthermore, Schick and Schütze [5] suggest that small models can be efficient few-shot learners, particularly when designed for narrow domains or specialised tasks. Observations from these studies are aligned with our findings, which demonstrate that several well-optimised small and mid-sized models achieve strong correctness while maintaining low memory usage. This supports the idea that well-targeted and task-adapted SLMs can offer practical performance advantages without the high memory demands typically associated with larger models.
@@ 4.3 Multilingual Capabilities and Limitations of SLMs (RQ3)
Performance Analysis of SLMs in Multilingual Code Generation: Our multilingual performance analysis shows that the performance of SLMs appears to fluctuate across programming languages; however, the analysis reveals that these variations are not statistically significant. This suggests that overall, SLMs maintain a reasonably consistent ability to generalise across multiple programming languages. Nevertheless, practical patterns are still observable as models tend to perform better in languages such as Python, Java, and PHP, while showing comparatively weaker results in Go, C++ and Ruby. The possible reasons for these trends include: (i) the uneven representation of programming languages in pretraining datasets, and (ii) variations in syntactic structure across languages.
Prior studies have shown that languages like C++ and Ruby introduce greater syntactic depth, semantic ambiguity, and structural irregularity, which pose challenges for code generation models to handle consistently. Raychev et al. [22] explained that C++ is a complex language because of features like templates and unclear behaviour in some situations, which can make it harder for models to generate correct code. Similarly, Allamanis et al. [23] noted that unusual grammar rules and language-specific features in languages like Ruby can make it more challenging for models to perform well when working with multiple programming languages. Wang et al. [24] also observed that performance on different languages in code generation tasks can be strongly influenced by how well those languages are represented in the training data and by the model’s ability to generalize across structurally diverse programming paradigms.
@@ 5.1 Construct Validity
Construct validity addresses whether our evaluation accurately captures the intended outcomes, in this case, the code generation capabilities of SLMs. In our evaluation, we selected three established benchmarks, HumanEval, MBPP, and Mercury to assess the code generation capabilities of SLMs. These benchmarks provide a strong basis to measure how well SLMs generate functionally correct code, especially using the pass@k metric. However, the tasks in these benchmarks are generally limited in scale and complexity, algorithmic in nature, and focused on the Python programming language. This means they may not fully represent real-world programming tasks that involve larger, multi-module projects, asynchronous workflows, integration of legacy code, or domain-specific APIs. While we extended the evaluation to multiple programming languages through HumanEvalPack and CodeXGLUE, the distribution of programming languages still favours widely used languages like Python and Java, which could introduce a bias in model performance and limit generalizability to other programming languages such as Rust, Go, or TypeScript. Additionally, the use of pass@k and BLEU primarily focuses on correctness and surface-level similarity, but does not capture deeper aspects of code quality, such as maintainability, performance efficiency, idiomatic style, or security. Moreover, our evaluation does not incorporate human-in-the-loop qualitative assessments (e.g., developer ratings of code readability or usability), nor does it triangulate with real-world software development data, such as pull requests or issues from industrial repositories. This may lead to an incomplete assessment of how well models perform in practical software engineering contexts. Additionally, the use of static benchmarks and automated metrics provides only a partial view of model capabilities, potentially overlooking aspects such as adaptability to evolving codebases or interactive code refinement workflows. Although these design choices ensured a controlled and reproducible evaluation framework, they also constrain the external applicability of our findings. Our benchmark suite is diverse and well-accepted, but it may lack the ability to fully capture the complete range of capabilities required for robust and production-grade code generation.
@@ 5.2 Internal Validity
Internal validity for this study refers to the accuracy with which we conducted our experiments and whether we can trust the resulting data. To ensure fair and consistent evaluation, we carefully designed our experiments with fixed decoding settings (e.g., temperature = 0.2, top-p = 0.95, max tokens = 512), a standard zero-shot prompting format, and automated evaluation tools. While we used the same decoding configuration across all models to maintain comparability, decoding parameters can significantly influence generation diversity and correctness, which introduces sensitivity that may affect result stability. We used a zero-shot prompting strategy for all tasks, meaning the models received no in-context examples. Although this setup reflects a generalisation test, it may underestimate model performance relative to few-shot prompting or instruction tuning, which are commonly used in practice. Our experiments were conducted on two different hardware platforms: an NVIDIA L4 GPU and an RTX 3090 GPU, each equipped with 24 GB of VRAM. Although similar in memory, architectural differences such as CUDA cores, memory bandwidth, or driver optimisations may have influenced inference latency or resource usage. To mitigate this, each model was assigned to a single platform, with models evenly distributed across both platforms, and we ensured that both environments used the same software stack, CUDA version, batch sizes, and decoding settings. We also fixed the random seed to reduce output variability. These steps help ensure that hardware differences and configuration sensitivity do not systematically affect the results, thereby improving the consistency and reproducibility of our findings.
@@ 5.3 External Validity
External validity refers to how well our results apply outside of our specific experimental setup. In our study, we only included models with a decoder-only architecture. All of them were open-source and had up to 10 billion parameters. They were released between 2022 and 2024. This means we did not include other types of models, such as encoder-only or encoder-decoder models. We also did not evaluate larger models or proprietary ones, such as GPT-4 or Gemini. As a result, our findings are limited in scope and may not be applicable to other types of models. This design choice was intended to ensure comparability across open-source SLMs under similar resource constraints, but it limits generalisation to other model classes. Additionally, the benchmarks used for evaluating SLMs covered a limited number of programming languages (e.g., Python, Java), excluding niche or domain-specific languages such as R, Kotlin, or Swift. This may bias our evaluation toward languages that are better supported by current pretraining corpora, while under-representing specialised domains such as data science or mobile development. This limits the extent to which we can apply our findings. In addition, we used zero-shot prompting across all tasks to test how well models generalise without seeing any examples. However, it may not reflect many real-world coding scenarios, where developers often provide examples, context, or iterative feedback. Consequently, models optimised for instruction-following or few-shot learning may appear weaker under zero-shot evaluation. As a result, our evaluation may not accurately reflect how well some models perform when provided with additional guidance. To mitigate this, we used a standardised zero-shot format across all models to ensure fair comparisons, but future work should include few-shot and interactive prompting setups to better approximate practical usage. Models that rely heavily on context may appear weaker in this setup. While zero-shot evaluation helps reveal a models raw ability, it may limit our understanding of its full potential in interactive settings.
@@ 5.4 Conclusion Validity
Conclusion validity describes how well our results support the drawn conclusions. In this study, we employed statistical methods, including ANOVA and Tukey’s HSD post-hoc test, to determine whether the differences in performance and efficiency across models and benchmarks were statistically significant. These tests come with certain assumptions. First, the data points should be independent, which was satisfied since each model was evaluated separately. Second, the results within each group should have similar variance, which we achieved by grouping models based on size and maintaining consistent conditions. Third, the data should follow a roughly normal distribution, which is reasonable in our case because we used averages across many tasks. Lastly, the model groups were not randomly selected but carefully chosen, which meets the fixed-effects requirement of ANOVA. Given how we set up the experiments, all these assumptions were reasonably met. However, a few factors could still impact the results, such as timeouts or out-of-memory errors, particularly when larger models are tested on tasks that require more memory. Such failures may result in incomplete results for some models on specific benchmarks, potentially introducing bias. To reduce this issue, we adjusted batch sizes and limited the number of token generations where needed. We also documented any failed runs and excluded them consistently across models to avoid skewing comparisons. Some efficiency metrics, like inference speed, might also be affected by slight differences in system load or hardware behaviour. To mitigate this, we ensured that each model was run under controlled and consistent system configurations, with dedicated runs to minimise interference from background processes. Even though some slight variations may still exist, we observed consistent trends, which suggests our conclusions are reliable and would likely hold under similar conditions.
@@ 6 Related Work
With the rapid advancements in Artificial Intelligence (AI), Language Models have emerged as powerful tools for code generation [25]. Although LLMs such as GPT-4 [26] and Gemini [27] have demonstrated impressive capabilities, they continue to face practical challenges. Xu and et al.[1] conducted a comprehensive survey on resource-efficient LLMs and highlighted that both training and inference require high-performance GPUs, substantial memory bandwidth, significant power, and specialised cooling systems, which limit their widespread adoption. Wang et al. [2] discussed the use of SLMs as a more practical alternative. They showed that SLMs can perform well on specific tasks while using less memory. The authors also explained that SLMs are easier to deploy and can work better in real-world settings where computing resources are limited. However, our current understanding of the consistency and effectiveness of SLMs in the context of code generation remains underexplored.
@@
Recently, many LLMs have been explicitly developed focusing on coding and have been widely applied to tasks such as program synthesis, code completion, and bug fixing [28]. Chen et al.[6] introduced Codex, a code-specific LLM trained in extensive public code repositories. They discussed its ability to convert natural language instructions into executable code and reported its strong performance on the HumanEval benchmark, highlighting its practical utility in tools such as GitHub Copilot. Rozière et al. [7] introduced an open source alternative, Code Llama, offering competitive results in benchmarks and support for extended input contexts. The authors reported that its performance on the HumanEval and MBPP benchmarks is on par with some of the strongest open models available, particularly in zero-shot settings. Furthermore, Li et al. [29] introduced AlphaCode, which was tested on competitive programming challenges. They emphasised its ability to produce diverse and creative solutions that aligned closely with human-level problem-solving strategies, achieving results comparable to those of skilled human participants.
@@
In response to growing concerns about the resource demands of LLMs, recent studies have focused on developing SLMs that are more efficient and accessible. Wang et al. [2] emphasized that SLMs can significantly reduce the computational overhead while still achieving competent performance on various tasks. Among early SLMs, GPT-Neo [30] was introduced as a lightweight alternative capable of general-purpose code generation, with the authors highlighting its ability to run on limited hardware while maintaining competitive performance. Similarly, Xu et al. [31] conducted a systematic evaluation of code-focused LMs and noted that smaller models like PolyCoder were able to perform well in certain code understanding and generation scenarios, despite having significantly fewer parameters than traditional LLMs. More recently, Lozhkov et al. [32] introduced StarCoder2 as a scalable open-source family of code models designed to balance performance and efficiency, and reported that its smaller variants demonstrated promising results in zero-shot code generation tasks.
@@
A variety of benchmarks have been developed to evaluate the performance of language models in code generation, each targeting different aspects of functionality and generalisation. Hendrycks et al. [33] introduced APPS, a dataset designed to assess problem-solving ability through programming challenges of varying difficulty. Chen et al. [6] proposed HumanEval, a widely used benchmark where models are tested on their ability to generate correct Python functions from descriptions written in natural language. To further assess real-world programming capabilities, Austin et al. [34] released MBPP, which focuses on Python code snippets and includes manually-written test cases to evaluate correctness. These datasets have become standard tools for assessing LLMs and SLMs alike. Muennighoff et al. [35] introduced HumanEvalPack, extending these benchmarks into a multilingual setting. The authors highlighted that models trained primarily on English often struggle with syntax and semantics in different programming languages, making multilingual benchmarks essential for more comprehensive evaluations. Du et al. [36] introduced Mercury, a benchmark specifically designed to measure code efficiency, including runtime and memory usage, in addition to correctness. The authors emphasised that efficiency remains underexplored in model evaluation, despite being crucial for practical deployment. Across these benchmarks, the most common evaluation metric is pass@k, which measures whether at least one of the top-k generated outputs passes the unit tests. The BLEU score is also occasionally used to measure surface-level similarity between the generated and reference outputs. However, several studies caution that it may not fully capture functional correctness [28]. Chen et al. [28] further noted that differences in evaluation protocols, such as varying sampling temperatures, prompt formats, or test coverage, make it difficult to directly compare model performance across studies. This highlights the ongoing need for standardised evaluation methods to enable more reliable and reproducible assessments of code generation models.
@@
While much progress has been made in developing and evaluating LLMs for programming tasks, empirical studies that assess their integration and effectiveness in real-world SE workflows are still few. Zhang et al. [25] conducted a comprehensive survey on the use of LLMs in SE, covering areas such as program synthesis, code summarisation, and defect prediction. The authors noted that while LLMs show great promise, their actual deployment in SE practice is limited by scalability, interpretability, and cost effectiveness. Despite these efforts, most empirical studies have focused on LLMs such as Codex, with limited attention to SLMs. Subramanian et al. [3] explicitly identified this gap, noting that while SLMs are increasingly being proposed as cost-effective alternatives, their effectiveness in software engineering contexts remains largely untested. Likewise, Lu et al. [8] observed that empirical validation of SLMs across various programming tasks is often lacking, particularly in controlled environments that reflect real-world SE needs. This points out a critical research gap: although SLMs can offer promising advantages in terms of resource efficiency and deployment feasibility, their practical applicability in SE remains underexplored and under-evaluated.
@@
Conclusive summary: In summary, while LLMs have been extensively studied for code generation, there is a lack of empirical evaluation of SLMs in this context. Prior works have emphasised the potential of SLMs [2, 3], but their consistency and robustness in code generation tasks remain underexplored [8]. This work addresses this gap by benchmarking a range of SLMs across widely used datasets and evaluation metrics. Unlike earlier studies that focus on general NLP tasks (e.g., [37]), we focus specifically on SLMs that perform code generation tasks. Our empirical study reveals their actual capabilities and limitations in resource-constrained real-world development environments. This approach enables a more informed understanding of where and how SLMs can be effectively applied in software engineering workflows.

### v4

@@ 4.2 Efficiency Trade-offs in SLM Deployment (RQ2)
Resource Consumption and Inference Speed vs. Performance of SLMs: Our analysis reveals that GPU memory usage (VRAM) increases significantly with model size. At the same time, inference speed does not show a corresponding statistical variation. We observe this behavior because all models in our study are SLMs (0.4B–10B parameters), where differences in latency are naturally smaller than in large LLMs (e.g., >30B parameters). In addition, all models were evaluated using identical decoding settings and on relatively small, single-function benchmark tasks, which further reduces variation in inference speed. This indicates that, rather than inference speed, VRAM is the constraint when deploying SLMs in practice. As models scale up, their memory demands grow higher, creating significant challenges for deployment in resource-constrained environments. However, this does not imply that larger models are always necessary to achieve high performance. While high-performing SLMs often belong to the larger parameter groups, several smaller and mid-sized models demonstrate a favourable balance between correctness and resource efficiency. These models offer compelling alternatives for use cases where hardware limitations prohibit the use of resource-intensive models.
Prior studies (e.g.,[Lu2024SmallLM],[schick-schutze-2021-just]) have similarly emphasized this dynamic. While inference speed can often be optimised through software and hardware tuning, memory capacity, especially VRAM, remains a primary constraint in deploying small language models. Recent findings by Ashkboos et al. [ashkboos2024computational] indicate that even SLMs with modest parameter counts face significant memory bottlenecks during both training and inference, limiting their usability in resource-constrained settings. Lu et al.[Lu2024SmallLM] also observed that many performance evaluations overlook memory bottlenecks, leading to unrealistic assessments of deployment feasibility. Moreover, Subramanian et al.[Subramanian2025survey] emphasised treating efficiency not merely as another dimension of scale, but as an important metric in empirical evaluations of SLMs, especially in applied software engineering contexts. Furthermore, Schick and Schütze [schick-schutze-2021-just] suggest that small models can be efficient few-shot learners, particularly when designed for narrow domains or specialised tasks. Observations from these studies are aligned with our findings, which demonstrate that several well-optimised small and mid-sized models achieve strong correctness while maintaining low memory usage. This supports the idea that well-targeted and task-adapted SLMs can offer practical performance advantages without the high memory demands typically associated with larger models.
@@ 4.3 Multilingual Capabilities and Limitations of SLMs (RQ3)
Performance Analysis of SLMs in Multilingual Code Generation: Our multilingual performance analysis shows that the performance of SLMs appears to fluctuate across programming languages; however, the analysis reveals that these variations are not statistically significant. This suggests that overall, SLMs maintain a reasonably consistent ability to generalise across multiple programming languages. However, the multilingual benchmarks used in this study cover only a limited set of widely used programming languages; as a result, this generalizability applies only to the languages included in our study and cannot be assumed to extend to all programming languages. Nevertheless, practical patterns are still observable as models tend to perform better in languages such as Python, Java, and PHP, while showing comparatively weaker results in Go, C++ and Ruby. The possible reasons for these trends include: (i) the uneven representation of programming languages in pretraining datasets, and (ii) variations in syntactic structure across languages.
Prior studies have shown that languages like C++ and Ruby introduce greater syntactic depth, semantic ambiguity, and structural irregularity, which pose challenges for code generation models to handle consistently. Raychev et al. [Raychev] explained that C++ is a complex language because of features like templates and unclear behaviour in some situations, which can make it harder for models to generate correct code. Similarly, Allamanis et al. [Allamanis] noted that unusual grammar rules and language-specific features in languages like Ruby can make it more challenging for models to perform well when working with multiple programming languages. Wang et al. [wang2021codet5] also observed that performance on different languages in code generation tasks can be strongly influenced by how well those languages are represented in the training data and by the model’s ability to generalize across structurally diverse programming paradigms.
@@ 5.1 Construct Validity
Construct validity addresses whether our evaluation accurately captures the intended outcomes; in this study, the code-generation capabilities of SLMs. Our evaluation uses three established benchmarks: HumanEval, MBPP, and Mercury. These benchmarks provide a controlled, reproducible method for assessing functional correctness using the pass@k metric. However, the tasks in these benchmarks are generally small-scale, algorithmic in nature, and primarily focused on Python. This means they represent only a limited portion of real-world software engineering, which commonly requires multi-module codebases, asynchronous workflows, domain-specific APIs, and interaction with legacy systems. This limitation narrows the scope of our evaluation. To partially mitigate this constraint, we complement these benchmarks with multilingual benchmark datasets to reduce language-specific bias in the evaluation, although this does not fully eliminate the underlying constraint.
To expand our study coverage beyond Python-only evaluation, we included HumanEvalPack and CodeXGLUE benchmarks for multilingual assessment. This strategy broadens our analysis to several programming languages. However, both benchmark suites are still dominated by high-resource languages such as Python, Java, JavaScript, PHP, C++, and Go. Languages with smaller or rapidly evolving ecosystems, such as Rust, Kotlin, Swift, and TypeScript, are largely absent. As a result, the multilingual generalization observed in our results applies only to the languages included in these datasets. To address this limitation, we interpret our multilingual findings within the scope of the supported languages and avoid extending claims to broader cross-language generalization.
Furthermore, the use of automated metrics defined by the benchmark suites enables reproducible comparisons across a large number of models. However, it influences construct validity. Pass@k measures functional correctness, while BLEU captures surface-level similarity rather than deeper semantic or stylistic qualities. These metrics do not account for aspects of code quality such as maintainability, readability, performance efficiency, idiomatic usage, or security. Additionally, our study does not include human judgments of code quality, qualitative assessments of readability or usability, or triangulation with real-world software development data such as pull requests, bug reports, or production code artifacts. We acknowledge that the current scope of our study may not capture several dimensions of code quality that are important to practitioners. This limitation points to an opportunity for future work to integrate richer evaluation methods, such as human code reviews, static analysis tools, security assessments, and project-level tasks to achieve a more holistic understanding of SLM capabilities.
An additional concern for construct validity arises from potential contamination by benchmarks. Models can have been trained on datasets that include tasks similar or identical to the benchmark used in this study. To address this concern, we examined the models’ training data. Although some model families (e.g., Qwen2.5-coder) explicitly report filtering out datasets such as HumanEval and MBPP during pretraining, the full pretraining data of many models are not well documented. As a result, we cannot rule out this risk. Such overlap could inflate performance for certain models. To mitigate this limitation, we follow standard benchmarking practices used in prior work and acknowledge that potential training data leakage remains a limitation of our evaluation.
Finally, our use of static benchmarks and fully automated evaluation procedures provides consistency and reproducibility but may not reflect dynamic or interactive development workflows. Modern software engineering often involves iterative refinement, debugging, code review, and adaptation to evolving codebases—capabilities that static benchmark tasks cannot fully capture. Although these design choices enabled rigorous comparisons across models, they inevitably constrain the construct space represented in this study. While our benchmark suite is well-accepted and diverse, we acknowledge that it does not encompass the full range of capabilities required for robust, production-grade code generation. This limitation cannot be fully addressed within the scope of this work, and future research should incorporate more interactive and context-rich evaluation settings to better approximate real-world development scenarios.
@@ 5.2 Internal Validity
Internal validity for this study refers to the accuracy with which we conducted our experiments and to the extent to which we can trust the resulting data. To ensure fair and consistent evaluation, we carefully designed our experiments with fixed decoding settings (e.g., temperature = 0.2, top-p = 0.95, max tokens = 512), a standard zero-shot prompting format, and automated evaluation tools. These decoding parameters follow standard practice in code-generation benchmarks and match widely used configurations (e.g., official HumanEval implementations and MBPP evaluation scripts [openai_humaneval_2021], [google_mbpp_2021]), as well as vendor recommendations that suggest low-temperature sampling for deterministic code generation (e.g., NVIDIA NeMo documentation [nvidia_nemo_2025]). A limitation of this approach is that decoding parameters can affect generation diversity and correctness, leading to performance variations across configurations. Although a full sensitivity analysis could provide a more detailed understanding of how decoding settings affect model behavior, conducting such an analysis across all models would require exploring a large multidimensional parameter space, substantially increasing computational cost and reducing comparability with prior benchmark studies. To maintain methodological consistency, we therefore applied the same decoding configuration across all models. We acknowledge this as a limitation of this study, and future research should systematically examine the robustness of SLM performance under different decoding parameters.
We used a zero-shot prompting strategy for all tasks, meaning the models received no in-context examples. Although this setup reflects a generalization, it may underestimate model performance relative to few-shot prompting or instruction tuning, which are commonly used in practice. We choose this strategy because prior work shows that few-shot prompting predominantly benefits instruction-tuned models rather than base models. As all models in our study are based on SLMs, incorporating few-shot examples would introduce methodological bias and reduce comparability. Nonetheless, we acknowledge that an evaluation relying exclusively on zero-shot prompting limits the ability to assess model performance in example-driven or interactive development workflows. As a result, this remains a limitation of our evaluation: future work should evaluate SLMs across different prompting settings. Besides, our experiments were conducted on two different hardware platforms: an NVIDIA L4 GPU and an RTX 3090 GPU, each equipped with 24 GB of VRAM. This ensures that the primary efficiency metric (VRAM usage) in our analysis remains hardware independent. Although similar in memory, architectural differences such as CUDA cores, memory bandwidth, or driver optimisations can influence inference latency. To mitigate this issue, each model was evaluated entirely on a single GPU type, and models were evenly distributed across both platforms. In addition, both hardware environments used identical software stacks (CUDA version, driver version, batch sizes, decoding settings), and we fixed random seeds to reduce variability. These precautions ensure that any hardware-related variability appears as random noise rather than systematic bias. Besides, we acknowledge that using a single GPU type would be ideal. Still, the steps taken help ensure that hardware differences do not meaningfully affect the reliability of our comparative results.
@@ 5.3 External Validity
External validity refers to the extent to which our results generalise beyond the specific experimental setup. In our study, we only included models with a decoder-only architecture. All of them were open-source and had up to 10 billion parameters. They were released between 2022 and 2024. This means we did not include other model types, such as encoder-only or encoder-decoder models. We also did not evaluate larger models or proprietary ones, such as GPT-4 or Gemini. As a result, our findings are limited in scope and may not apply to other types of models. This design choice was intended to ensure comparability across open-source SLMs under similar resource constraints, but it limits generalisation to other model classes.
Another threat to external validity arises from the nature of the tasks used in this study. The selected benchmarks consist primarily of short, self-contained programming problems that do not capture key characteristics of real-world software engineering, such as multi-file project structures, long-range dependencies, integration with external libraries, or iterative development workflows. While these benchmarks are widely used in code-generation research and provide a controlled, reproducible evaluation environment, the evaluation results represent only a partial view of practical software development. We address this limitation in part by including multiple benchmark spanning different task types and languages; however, static benchmarks cannot fully reflect the complexities of production environments, where broader contextual understanding and architectural reasoning are required. We therefore acknowledge this as a limitation of the present work, and future research should incorporate project-level datasets or tasks drawn from real-world repositories to assess SLM behavior in practical settings more comprehensively. Moreover, the benchmarks used in this study evaluate SLMs on only a subset of languages and exclude niche or domain-specific languages common in specialized software engineering contexts. We partially address this limitation by incorporating multilingual benchmark suites, such as HumanEvalPack and CodeXGLUE, which extend coverage beyond Python and enable evaluation across multiple widely used languages. However, these benchmarks still represent only a limited portion of the broader programming language ecosystem. As a result, our multilingual findings should be interpreted within the scope of the languages included in these datasets and may not generalize to specialized or domain-specific languages. We acknowledge this limitation, and future work should evaluate SLMs on more diverse, domain-relevant language sets to improve generalizability
A further limitation concerns the efficiency dimensions evaluated in this study. Our assessment of computational efficiency focused on VRAM usage and inference time, as these metrics are commonly reported in prior work and can be measured reliably in GPU-based environments. However, this approach does not capture other important aspects of efficiency, such as energy consumption, CPU utilization, thermal characteristics, or power efficiency. These factors are particularly relevant for edge devices and low-power deployment scenarios [s23031279]. These dimensions were outside the scope of this work. As a result, our findings represent only a partial view of the practical deployment considerations for SLMs. We acknowledge this limitation and identify the inclusion of energy profiling and hardware-level performance measurements as an important direction for future research to provide a more comprehensive understanding of real-world deployment.
@@
In addition, we used zero-shot prompting across all tasks to test how well models generalise without seeing any examples. However, it may not reflect many real-world coding scenarios, where developers often provide examples, context, or iterative feedback. Consequently, models optimised for instruction-following or few-shot learning may appear weaker under zero-shot evaluation. As a result, our evaluation may not accurately reflect how well some models perform when provided with additional guidance. To mitigate this, we used a standardised zero-shot format across all models to ensure fair comparisons, but future work should include few-shot and interactive prompting setups to approximate practical usage better. Models that rely heavily on context may appear weaker in this setup. While zero-shot evaluation helps reveal a models raw ability, it may limit our understanding of its full potential in interactive settings.
@@ 5.4 Conclusion Validity
Conclusion validity describes how well our results support the drawn conclusions. In this study, we employed statistical methods, including ANOVA and Tukey’s HSD post hoc test, to assess whether differences in performance and efficiency across models and benchmarks were statistically significant. These tests come with certain assumptions. First, the data points should be independent, which was satisfied since each model was evaluated separately. Second, the results within each group should exhibit similar variance, which we achieved by grouping models by size and maintaining consistent conditions. Third, the data should follow a roughly normal distribution, which is reasonable in our case because we used averages across many tasks. Lastly, the model groups were not randomly selected but carefully chosen, which meets the fixed-effects requirement of ANOVA. Given how we set up the experiments, all these assumptions were reasonably met. However, several factors may still affect the results, such as timeouts or out-of-memory errors, particularly when larger models are tested on tasks that require more memory. Such failures may result in incomplete results for some models on specific benchmarks, potentially introducing bias. To reduce this issue, we adjusted batch sizes and limited the number of token generations where needed. We also documented any failed runs and excluded them consistently across models to avoid skewing comparisons. Some efficiency metrics, such as inference speed, may also be affected by slight differences in system load or hardware behavior. To mitigate this, we ensured that each model was run under controlled and consistent system configurations, with dedicated runs to minimize interference from background processes. Although some slight variations may still exist, we observed consistent trends, suggesting that our conclusions are reliable and would likely hold under similar conditions.
@@ 6 Related Work
With the rapid advancements in Artificial Intelligence (AI), Language Models have emerged as powerful tools for code generation [zhang2023survey]. Although LLMs such as GPT-4 [GPT-4] and Gemini [team2023gemini] have demonstrated impressive capabilities, they continue to face practical challenges. Xu et al.[xu2024survey] conducted a comprehensive survey on resource-efficient LLMs and highlighted that both training and inference require high-performance GPUs, substantial memory bandwidth, significant power, and specialised cooling systems, which limit their widespread adoption. Wang et al. [Wang2024ACS] discussed the use of SLMs as a more practical alternative. They showed that SLMs can perform well on specific tasks while using less memory. The authors also explained that SLMs are easier to deploy and can work better in real-world settings where computing resources are limited. However, our current understanding of the consistency and effectiveness of SLMs in the context of code generation remains underexplored.
@@
Recently, many LLMs have been explicitly developed focusing on coding and have been widely applied to tasks such as program synthesis, code completion, and bug fixing [chen2024survey]. Chen et al.[Chen2021EvaluatingLL] introduced Codex, a code-specific LLM trained in extensive public code repositories. They discussed its ability to convert natural language instructions into executable code and reported its strong performance on the HumanEval benchmark, highlighting its practical utility in tools such as GitHub Copilot. Rozière et al. [CodeLlama] introduced an open source alternative, Code Llama, offering competitive results in benchmarks and support for extended input contexts. The authors reported that its performance on the HumanEval and MBPP benchmarks is on par with some of the strongest open models available, particularly in zero-shot settings. Furthermore, Li et al. [Li@2022] introduced AlphaCode, which was tested on competitive programming challenges. They emphasised its ability to produce diverse and creative solutions that aligned closely with human-level problem-solving strategies, achieving results comparable to those of skilled human participants.
@@
In response to growing concerns about the resource demands of LLMs, recent studies have focused on developing SLMs that are more efficient and accessible. Wang et al. [Wang2024ACS] emphasized that SLMs can significantly reduce the computational overhead while still achieving competent performance on various tasks. Among early SLMs, GPT-Neo [gpt-neo] was introduced as a lightweight alternative capable of general-purpose code generation, with the authors highlighting its ability to run on limited hardware while maintaining competitive performance. Similarly, Xu et al. [xu2022systematic] conducted a systematic evaluation of code-focused LMs and noted that smaller models like PolyCoder were able to perform well in certain code understanding and generation scenarios, despite having significantly fewer parameters than traditional LLMs. More recently, Lozhkov et al. [lozhkov2024starcoder] introduced StarCoder2 as a scalable open-source family of code models designed to balance performance and efficiency, and reported that its smaller variants demonstrated promising results in zero-shot code generation tasks.
@@
A variety of benchmarks have been developed to evaluate the performance of language models in code generation, each targeting different aspects of functionality and generalisation. Hendrycks et al. [hendrycks2021measuring] introduced APPS, a dataset designed to assess problem-solving ability through programming challenges of varying difficulty. Chen et al. [Chen2021EvaluatingLL] proposed HumanEval, a widely used benchmark where models are tested on their ability to generate correct Python functions from descriptions written in natural language. To further assess real-world programming capabilities, Austin et al. [austin2021program] released MBPP, which focuses on Python code snippets and includes manually-written test cases to evaluate correctness. These datasets have become standard tools for assessing LLMs and SLMs alike. Muennighoff et al. [OctoPack] introduced HumanEvalPack, extending these benchmarks into a multilingual setting. The authors highlighted that models trained primarily on English often struggle with syntax and semantics in different programming languages, making multilingual benchmarks essential for more comprehensive evaluations. Du et al. [Du2024MercuryAC] introduced Mercury, a benchmark specifically designed to measure code efficiency, including runtime and memory usage, in addition to correctness. The authors emphasised that efficiency remains underexplored in model evaluation, despite being crucial for practical deployment. Across these benchmarks, the most common evaluation metric is pass@k, which measures whether at least one of the top-k generated outputs passes the unit tests. The BLEU score is also occasionally used to measure surface-level similarity between the generated and reference outputs. However, several studies caution that it may not fully capture functional correctness [chen2024survey]. Chen et al. [chen2024survey] further noted that differences in evaluation protocols, such as varying sampling temperatures, prompt formats, or test coverage, make it difficult to directly compare model performance across studies. This highlights the ongoing need for standardised evaluation methods to enable more reliable and reproducible assessments of code generation models.
@@
While much progress has been made in developing and evaluating LLMs for programming tasks, empirical studies that assess their integration and effectiveness in real-world SE workflows are still few. Zhang et al. [zhang2023survey] conducted a comprehensive survey on the use of LLMs in SE, covering areas such as program synthesis, code summarisation, and defect prediction. The authors noted that while LLMs show great promise, their actual deployment in SE practice is limited by scalability, interpretability, and cost effectiveness. Despite these efforts, most empirical studies have focused on LLMs such as Codex, with limited attention to SLMs. Subramanian et al. [Subramanian2025survey] explicitly identified this gap, noting that while SLMs are increasingly being proposed as cost-effective alternatives, their effectiveness in software engineering contexts remains largely untested. Likewise, Lu et al. [Lu2024SmallLM] observed that empirical validation of SLMs across various programming tasks is often lacking, particularly in controlled environments that reflect real-world SE needs. This points out a critical research gap: although SLMs can offer promising advantages in terms of resource efficiency and deployment feasibility, their practical applicability in SE remains underexplored and under-evaluated.
@@
Conclusive summary: In summary, while LLMs have been extensively studied for code generation, there is a lack of empirical evaluation of SLMs in this context. Prior works have emphasised the potential of SLMs [Wang2024ACS, Subramanian2025survey], but their consistency and robustness in code generation tasks remain underexplored [Lu2024SmallLM]. This work addresses this gap by benchmarking a range of SLMs across widely used datasets and evaluation metrics. Unlike earlier studies that focus on general NLP tasks (e.g., [Lepagnol2024SmallLM]), we focus specifically on SLMs that perform code generation tasks. Our empirical study reveals their actual capabilities and limitations in resource-constrained real-world development environments. This approach enables a more informed understanding of where and how SLMs can be effectively applied in software engineering workflows.